# Tropical Gradient Descent

### v2

@@ 2.3 Tropical Data Science
While young, the field of tropical data science is motivated by the crucial question of statistical learning from the increasing volume of phylogenetic tree data (Kahn,, 2011). In contrast to the BHV (Billera et al.,, 2001) or Robinson–Foulds metrics (Robinson & Foulds,, 1981), the tropical interpretation of tree space provides both interpretable geometry and computational efficiency (Steel & Penny,, 1993; Monod et al.,, 2018). Many of the statistical methods defined for the tropical setting are framed as optimisation problems (Barnhill et al.,, 2024b), and in this subsection we review some such tropical statistical optimization problems which we use to test our tropical descent methodology. We note the varying degrees of convexity demonstrated by each of them. While we restrict our considerations through the rest of this paper to statistics which arise from tropical location problems, in Appendix A we discuss other tropical statistical optimization problems.
@@ 2.3.1 Centrality Statistics
Our first problems of interest, Fermat–Weber points and Fréchet means, are centrality statistics for data on a general metric space; Fermat–Weber points (Drezner et al.,, 2004) act as a generalized median, while Fréchet means (Fréchet,, 1948) are a generalization of the mean.
Definition 19 (Tropical Fermat–Weber Points (Lin & Yoshida,, 2018)).
@@
Definition 20 (Tropical Fréchet means (Monod et al.,, 2018)).
@@ 2.3.2 Wasserstein Projections
The next statistical optimization problem was motivated by Cai & Lim, (2022), and looks to identify a tropical projection which minimizes the Wasserstein distance between samples in different tropical projective tori.
Definition 21 (Tropical Wasserstein Projections (Talbut et al.,, 2023)).
@@
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{p}(\mathbf{t})=\left(\frac{1}{K}\sum_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}^{p}\right)^{1/p}$
@@
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{\infty}(\mathbf{t})=\max_{k\leq K}\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{k}\|_{\mathrm{tr}}$
In contrast to the work by Lee et al., (2022) which looks to solve the optimal transport problem for intrinsic measures on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, this $p$-Wasserstein projection problem is computing an optimal projection for empirical samples.
@@
The tropical $p,\infty$-Wasserstein projections are min-tropical location problems, and the $\infty$-Wasserstein projection has $\min$-tropically convex sublevel sets.
@@
$\|[\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i}]_{j}-\mathbf{y}_{j}\|_{\mathrm{tr}}$ $=\max_{b}\max_{a\in J_{b}}[z_{ka}-t_{a}]-\min_{j}\max_{i\in J_{j}}[z_{ki}-t_{i}]$
$=\max_{a}[z_{ka}-t_{a}]+\max_{j}\min_{i\in J_{j}}[t_{i}-z_{ki}]$
$=\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]$
@@
$f_{p}(\mathbf{t})$ $=\left(\frac{1}{K}\sum_{k\leq K}\left(\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]\right)^{p}\right)^{1/p}$
$f_{\infty}(\mathbf{t})$ $=\max_{k\leq K}\max_{j}[\min_{i\in J_{j}}[t_{i}-z_{ki}-\min_{a}[t_{a}-z_{ka}]]]$
@@
As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sublevel sets are min-tropically convex by Proposition 15. This is preserved when taking a maximum over samples, so $f_{\infty}$ also has $\min$-tropically convex sublevel sets. ∎
@@ 3 Steepest Descent with the Tropical Norm
We treat optimization over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as optimization over $\mathbb{R}^{N}$ with the knowledge that $f$ is constant along $\mathbf{1}$-rays. We assume we can compute some gradient-like function $\nabla f:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ which is satisfying Assumption 23.
@@
Definition 24 (Generalized Clarke Derivative (2.13, I.I of Kiwiel, (1985))).
@@
Proof Suppose that $\nabla f(\mathbf{t})_{j}<0$. Then there is a sequence of differentiable $\mathbf{y}_{\ell}$ satisfying $\nabla f(\mathbf{y}_{\ell})_{j}<0$, $\mathbf{y}_{\ell}\rightarrow\mathbf{t}$. Hence by the chain rule, as $g$ is increasing in every $h_{k}$, there is some $h_{k}$ which is strictly decreasing in the $j^{th}$ coordinate at $\mathbf{y}_{\ell}$. WLOG, the same $h_{k}$ is strictly decreasing in the $j^{th}$ coordinate for all $\mathbf{y}_{\ell}$. As $\gamma$ is increasing in every coordinate, $h_{k}=\hat{\gamma}(\mathbf{y}_{\ell}-\mathbf{x}_{k})$ is strictly decreasing in the $j^{th}$ coordinate if and only if $j\in\text{argmin}_{i}(y_{\ell i}-x_{ki})$. This is preserved through the limit as $\ell\rightarrow\infty$, so we conclude that $j\in\text{argmin}_{i}[t_{i}-x_{i}]$. ∎
In practice, rather than computing elements of the generalized Clarke derivative, we use PyTorch automatic differentiation (Paszke et al.,, 2019). The derivatives computed by PyTorch are dependent on the computational representation of a function, but are designed to respect the chain rule. We must therefore ensure that our computational representation of $g$ has non-negative derivatives in each $h_{k}$, while each $h_{k}$ is computed such that its $t_{j}$ derivative is negative only if $i\in\text{argmin}_{i}(t_{i}-x_{ki})$. For such a computational representation of $f$, Assumption 23 will hold by the chain rule. For all the tropical location problems we consider in this paper, we find their most natural computational representation gives rise to derivatives $\nabla f$ satisfying Assumption 23.
@@ 3.1 Proposed Method
Definition 26 (Steepest Descent Direction (Boyd & Vandenberghe,, 2004)).
[…]
@@
With respect to the Euclidean norm, the normalized derivative is a steepest descent direction. When we use a the tropical norm, the steepest descent directions are characterized by the following lemma.
@@
Suppose $\nabla f\neq\mathbf{0}$. Then the direction $\mathbf{v}$ is a normalized tropical steepest descent direction iff
$v_{i}=1$ $\Leftrightarrow\nabla f_{i}<0$
$v_{i}=0$ $\Leftrightarrow\nabla f_{i}>0$
@@
Throughout the rest of this paper, we use the min-tropical descent direction as convention, and will specify when the max-tropical descent direction is used instead. Tropical descent therefore refers to the method outlined by Algorithm 1, for some specified step size sequence $(a_{n})_{n\geq 1}$, while we refer to classical gradient descent as classical descent.
@@
for $n<$ max_steps do
     $\mathbf{g}_{ni}\leftarrow\nabla f_{i}(\mathbf{t}_{n-1})$;
     $\mathbf{d}_{ni}\leftarrow\mathbf{1}_{\mathbf{g}_{ni}<0}$;
     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n-1}+a_{n}\mathbf{d}_{n}$;
We note that over a large number of steps, Algorithm 1 can produce large $t_{ni}$ values as each coordinate is non-decreasing in $n$. If this results in overflow or precision errors, a normalisation operation can be included in each step, such as subtracting $\frac{1}{N}\sum_{i}t_{ni}$ from each $t_{ni}$. The function $f$ is invariant under such a translation due to the equivalence relation on the tropical projective torus.
In (Barnhill et al.,, 2024a, b), classical descent is implemented for the computation of Fermat–Weber points; this is the most natural choice as the Fermat–Weber problem is classically convex. However, the linear regression and Wasserstein projection problems lack classical convexity, and hence do not consistently produce accurate solutions via classical descent. It is these tropically convex problems for which tropical descent is necessary.
@@ 3.2 Convergence Guarantees
In this section we prove the theoretical guarantees of tropical descent for min-tropical location problems. We show that tropical descent must converge to the tropical convex hull of the data, and hence will necessarily find a minimum of a $\triangle_{\min}$-star-convex problem. Throughout this section, we assume the step sizes $a_{n}$ satisfy $a_{n}\rightarrow 0,\sum a_{n}=\infty$.
We first highlight the key difference in stability between classical descent and tropical descent; if some component of the derivative is locally zero, tropical descent cannot converge in that neighbourhood.
@@
Let $\mathbf{t}_{n}$ be some sequence defined by tropical descent. Assume that for all $n$, $\nabla f(\mathbf{t}_{n})\neq\mathbf{0}$. Suppose in some bounded open $U$, $\nabla f_{i}$ is identically 0. Then no point in $\operatorname{int}U$ is a convergence point for $\mathbf{t}_{n}$; in particular, if $\mathbf{t}_{m}\in U$ there is some $n>m$ such that $\mathbf{t}_{n}\notin U$.
@@
$\sum_{j}t_{nj}\geq\sum_{j}t_{mj}+\sum_{n\geq m}a_{n}\rightarrow\infty.$
@@
Suppose $f$ is a min-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Let $V=\operatorname{tconv}_{\max}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta_{j}(\mathbf{t})=\sum_{k\in[K]}\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
For any $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}<\epsilon/N$. Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$, where $s_{m}$ is the partial sum of the step size sequence $a_{m}$. Then for all $m\geq M_{2}$ we have:
$d_{{\rm tr}}(\mathbf{t}_{n},V)\leq\epsilon.$
@@
$\delta_{j}(\mathbf{t})=\min_{k\in[K]}\left[\sum_{i\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}\right].$
@@
Proof As discussed by Maclagan & Sturmfels, (2015), there is a well-defined metric projection map $\pi_{V}$ onto $V$ given by:
@@
[…]
@@
Proof For all $k\in[K],j\in[N]$, we have that $t_{mj}-x_{kj}-t_{mi}+x_{ki}$ is $1$-Lipschitz in $\mathbf{t}$ (with respect to $d_{\mathrm{tr}}$). Therefore $\delta_{j}$ is $N$-Lipschitz in $\mathbf{t}$. By the tropical descent rule:
$\delta_{j}(t_{m+1})$ $\leq\delta_{j}(\mathbf{t}_{m})+Nd_{\mathrm{tr}}(\mathbf{t}_{m},\mathbf{t}_{m+1})$
@@
We have assumed that for each $m$, $\nabla f\neq\mathbf{0}$ and so there is some $i$ such that $\nabla f_{i}<0$, and by Lemma 25, there is some $k$ such that $i\in\text{argmin}_{\ell}t_{m\ell}-v_{k\ell}$. Hence:
$\forall\,\ell:\quad t_{mj}-v_{kj}-t_{mi}+v_{ki}$ $\geq t_{mj}-v_{kj}-t_{m\ell}+v_{k\ell},$
$\Rightarrow\quad t_{mj}-v_{kj}-t_{mi}+v_{ki}$ $\geq\delta_{j}(\mathbf{t}_{m})/N,$
@@
Finally, from the claims above, we conclude that $\forall m\geq M_{2},i\in[N]$, we have $\delta_{i}(\mathbf{t}_{m})<\epsilon$, and so by our first claim, we have $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon$. ∎
As proven by Comăneci, (2024), tropical location problems have minima in the tropical convex hull of the data. The theorem above says that according to tropical descent, only such minima can be stable. We cannot do much better in terms of convergence results for general location problems; they can have disconnected sublevel sets so we would not expect gradient methods to find a global minimum. We would require some degree of convexity to guarantee global solvability, which is exactly what the following corollary gives us; tropical descent will necessarily converge to a minimum of any $\triangle_{\min}$-star-quasi-convex function.
@@
Let $f(\mathbf{t})$ be a $\triangle_{\min}$-star-quasi-convex function, and let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined by tropical descent such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Then there is some global minimum $\mathbf{t}^{*}$ such that for all $m\leq M_{2}$, $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\epsilon$ where $M_{2}$ is as in Theorem 29.
@@
Suppose $f$ is a max-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{n})_{n\geq 1}$ be a sequence of points defined via max-tropical descent directions, such that for all $n$: $\nabla f(\mathbf{t}_{n})\neq 0$. Let $V=\operatorname{tconv}_{\min}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta^{\prime}_{i}(\mathbf{t})=\sum_{k\in[K]}\sum_{j\in[N]}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
@@
The results above indicate convergence to $V$ or $\mathbf{t}^{*}$ in the tropical topology of $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, but $\mathbf{t}_{n}$ cannot not converge in the Euclidean topology of $\mathbb{R}^{N}$ due to the divergent partial sums of $\alpha_{n}$.
We conclude by establishing the convergence rate of tropical descent in practice; we take step sizes given by $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$, and assume derivatives have bounded tropical norm as is the case for piecewise linear 1-Lipschitz objective functions $h_{k}$. In this case, we prove convergence at a rate of $O(\sqrt{m})$.
@@
Suppose $f$ is a min-tropical location problem as in Theorem 29. Assume that there is some $L\in\mathbb{N}$ such that for all $\mathbf{t}_{m}$, $1/L\leq\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}\leq 2$, and assume further there is some $D$ such that $\forall k\leq K,n\in\mathbb{N}:\,d_{{\rm tr}}(\mathbf{t}_{n},\mathbf{x}_{k})\leq D$. We define $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$. Then for all $m$ satisfying
$1<\xi(m)\coloneqq\frac{\sqrt{m+1}}{2}-\frac{KDNL(N-1)}{8\alpha}+\sqrt{\left(\frac{\sqrt{m+1}}{2}-\frac{KDNL(N-1)}{8\alpha}\right)^{2}+LKN-1},$
@@
$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\frac{2\alpha N}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}.$

### v3

@@ 2.3 Tropical Data Science
While young, the field of tropical data science is motivated by the crucial question of statistical learning from the increasing volume of phylogenetic tree data (kahn2011future). In contrast to the BHV (billera2001geometry) or Robinson–Foulds metrics (ROBINSON1981131), the tropical interpretation of tree space provides both interpretable geometry and computational efficiency (penny1993distributions; monod2018StatPerspective). Many of the statistical methods defined for the tropical setting are framed as optimization problems (barnhill2023tml), and in this subsection we review some such tropical statistical optimization problems which we use to test our tropical descent methodology. We note the varying degrees of convexity demonstrated by each of them. While we restrict our considerations through the rest of this paper to statistics which arise from tropical location problems, in Appendix A we discuss other tropical statistical optimization problems.
@@ 2.3.1 Centrality Statistics
Our first problems of interest, Fermat–Weber points and Fréchet means, are centrality statistics for data on a general metric space; Fermat–Weber points (wesolowsky1993weber) act as a generalized median, while Fréchet means (frechet1948elements) are a generalization of the mean.
Definition 19 (Tropical Fermat–Weber Points (lin2018FermatWeber)).
@@
Definition 20 (Tropical Fréchet means (monod2018StatPerspective)).
@@ 2.3.2 Wasserstein Projections
The next statistical optimization problem was motivated by cai2022distances, and looks to identify a tropical projection which minimizes the Wasserstein distance between samples in different tropical projective tori.
Definition 21 (Tropical Wasserstein Projections (talbut2023probability)).
@@
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{p}(\mathbf{t})=\left(\frac{1}{K}\sum_{k\leq K}\|(\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i})_{{j\leq M}}-\mathbf{y}_{k}\|_{\mathrm{tr}}^{p}\right)^{1/p}$
@@
$\min_{\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}}f_{\infty}(\mathbf{t})=\max_{k\leq K}\|(\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i})_{{j\leq M}}-\mathbf{y}_{k}\|_{\mathrm{tr}}$
In contrast to the work by lee2021tropical which looks to solve the optimal transport problem for intrinsic measures on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, this $p$-Wasserstein projection problem is computing an optimal projection for empirical samples.
@@
The tropical $p,\infty$-Wasserstein projections are min-tropical location problems, and the $\infty$-Wasserstein projection has $\min$-tropically convex sub-level sets.
@@
$\|(\max_{i\in J_{j}}(\mathbf{x}_{k}-\mathbf{t})_{i})_{{j\leq M}}-\mathbf{y}_{k}\|_{\mathrm{tr}}$ $=\max_{b}\max_{a\in J_{b}}(z_{ka}-t_{a})-\min_{j}\max_{i\in J_{j}}(z_{ki}-t_{i})$
$=\max_{a}(z_{ka}-t_{a})+\max_{j}\min_{i\in J_{j}}(t_{i}-z_{ki})$
$=\max_{j}(\min_{i\in J_{j}}(t_{i}-z_{ki}-\min_{a}(t_{a}-z_{ka}{)))}$
@@
$f_{p}(\mathbf{t})$ $=\left(\frac{1}{K}\sum_{k\leq K}\left(\max_{j}(\min_{i\in J_{j}}(t_{i}-z_{ki}-\min_{a}(t_{a}-z_{ka}{)))}\right)^{p}\right)^{1/p}$
$f_{\infty}(\mathbf{t})$ $=\max_{k\leq K}\max_{j}(\min_{i\in J_{j}}(t_{i}-z_{ki}-\min_{a}(t_{a}-z_{ka}{)))}$
@@
As the $h_{k}$ are $\triangle_{\min}$-star-convex, their sub-level sets are min-tropically convex by Proposition 15. This is preserved when taking a maximum over samples, so $f_{\infty}$ also has $\min$-tropically convex sub-level sets. ∎
@@ 3 Steepest Descent with the Tropical Norm
We treat optimization over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ as optimization over $\mathbb{R}^{N}$ with the knowledge that $f$ is constant along $\mathbf{1}$-rays. We assume we can compute some gradient-like function $\nabla f:\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$ which satisfies Assumption 23.
@@
Definition 24 (Generalized Clarke Derivative (2.13, I.I of kiwiel2006methods)).
@@
Proof Suppose that $\nabla f(\mathbf{t})_{j}<0$. Then there is a sequence of differentiable $\mathbf{y}_{\ell}$ satisfying $\nabla f(\mathbf{y}_{\ell})_{j}<0$, $\mathbf{y}_{\ell}\rightarrow\mathbf{t}$. Hence by the chain rule, as $g$ is increasing in every $h_{k}$, there is some $h_{k}$ which is strictly decreasing in the $j^{th}$ coordinate at $\mathbf{y}_{\ell}$. WLOG, the same $h_{k}$ is strictly decreasing in the $j^{th}$ coordinate for all $\mathbf{y}_{\ell}$. As $\gamma$ is increasing in every coordinate, $h_{k}=\hat{\gamma}(\mathbf{y}_{\ell}-\mathbf{x}_{k})$ is strictly decreasing in the $j^{th}$ coordinate if and only if $j\in\text{argmin}_{i}(y_{\ell i}-x_{ki})$. This is preserved through the limit as $\ell\rightarrow\infty$, so we conclude that $j\in\text{argmin}_{i}(t_{i}-x_{i})$. ∎
In practice, rather than computing elements of the generalized Clarke derivative, we use PyTorch automatic differentiation (paszke2019pytorch). The derivatives computed by PyTorch are dependent on the computational representation of a function, but are designed to respect the chain rule. We must therefore ensure that our computational representation of $g$ has non-negative derivatives in each $h_{k}$, while each $h_{k}$ is computed such that its $t_{j}$ derivative is negative only if $j\in\text{argmin}_{i}(t_{i}-x_{ki})$. For such a computational representation of $f$, Assumption 23 will hold by the chain rule. For all the tropical location problems we consider in this paper, we find their most natural computational representation gives rise to derivatives $\nabla f$ satisfying Assumption 23.
@@ 3.1 Proposed Method
Definition 26 (Steepest Descent Direction (boyd2004convex)).
A normalized steepest descent direction at $\mathbf{t}$ with respect to the norm $\|\cdot\|$ is:
@@
With respect to the Euclidean norm, the normalized derivative is a steepest descent direction. The steepest descent directions under the tropical norm are characterized by the following lemma.
@@
Suppose $\nabla f\neq\mathbf{0}$. Then the direction $\mathbf{v}$ is a normalized tropical steepest descent direction (up to $\sim$ equivalence) iff
$\nabla f_{i}<0$ $\Rightarrow v_{i}=1,$
$\nabla f_{i}>0$ $\Rightarrow v_{i}=0.$
@@
Throughout the rest of this paper, we use the min-tropical descent direction as convention, and will specify when the max-tropical descent direction is used instead. Tropical descent therefore refers to the method outlined by Algorithm 1, for some specified step size sequence $(a_{m})_{m\geq 1}$, while we refer to classical gradient descent as classical descent.
@@
for $m<$ max_steps do
  $\mathbf{g}_{mi}\leftarrow\nabla f(\mathbf{t}_{m-1})_{i}$;
  $\mathbf{d}_{mi}\leftarrow\mathbf{1}_{\mathbf{g}_{mi}<0}$;
  $\mathbf{t}_{m}\leftarrow\mathbf{t}_{m-1}+a_{m}\mathbf{d}_{m}$;
We note that over a large number of steps, Algorithm 1 can produce large $t_{mi}$ values as each coordinate is non-decreasing in $m$. If this results in overflow or precision errors, a normalization operation can be included in each step, such as subtracting $\frac{1}{N}\sum_{{i\leq N}}t_{mi}$ from each $t_{mi}$. The function $f$ is invariant under such a translation due to the equivalence relation on the tropical projective torus.
We note that barnhill2024polytropes and barnhill2023tml implement classical descent for the computation of Fermat–Weber points; this is the most natural choice as the Fermat–Weber problem is classically convex. However, the linear regression and Wasserstein projection problems lack classical convexity, and hence do not consistently produce accurate solutions via classical descent. It is these tropically convex problems for which tropical descent is necessary.
@@ 3.2 Convergence Guarantees
In this section we prove the theoretical guarantees of tropical descent for min-tropical location problems. We show that tropical descent must converge to the tropical convex hull of the data, and hence will necessarily find a minimum of a $\triangle_{\min}$-star-convex problem. Throughout this section, we assume the step sizes $a_{m}$ satisfy $a_{m}\rightarrow 0,\sum a_{m}=\infty$.
We first highlight the key difference in stability between classical descent and tropical descent; if some component of the derivative is locally zero, tropical descent cannot converge in that neighborhood.
@@
Let $\mathbf{t}_{m}$ be some sequence defined by tropical descent. Assume that for all $m$, $\nabla f(\mathbf{t}_{m})\neq\mathbf{0}$. Suppose in some bounded open $U$, $\nabla f_{i}$ is identically 0. Then no point in $U$ is a convergence point for $\mathbf{t}_{m}$; in particular, if $\mathbf{t}_{m}\in U$ there is some $n>m$ such that $\mathbf{t}_{n}\notin U$.
@@
$\sum_{j}t_{nj}\geq\sum_{j}t_{mj}+\sum_{n\geq m^{\prime}\geq m}a_{m^{\prime}}\rightarrow\infty.$
@@
Suppose $f$ is a min-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{m})_{m\geq 1}$ be a sequence of points defined by tropical descent, such that for all $m$: $\nabla f(\mathbf{t}_{m})\neq 0$. Let $V=\operatorname{tconv}_{\max}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta_{j}(\mathbf{t})=\sum_{{k\leq K}}\sum_{{i\leq N}}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
For any $\epsilon>0$, let $M_{1}$ be such that for all $m\geq M_{1}$, $a_{m}<\epsilon/N$. Let $M_{2}$ be such that for all $m\geq M_{2}$ $s_{m}\geq\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$, where $s_{m}$ is the partial sum of the step size sequence $a_{m}$. Then for all $m\geq M_{2}$ we have:
$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon.$
@@
$\delta_{j}(\mathbf{t})=\min_{{k\leq K}}\left[\sum_{{i\leq N}}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}\right].$
@@
Proof As discussed by maclagan2015introduction, there is a well-defined metric projection map $\pi_{V}$ onto $V$ given by:
@@
 $\text{ where }\lambda_{k}=\max\{\lambda\in\mathbb{R}:\,\lambda\odot\mathbf{x}_{k}\boxplus\mathbf{t}=\mathbf{t}\}=\min_{i}(t_{i}-x_{ki}).$
@@
Proof For all $k\leq K,j\leq N$}, we have that $t_{mj}-x_{kj}-t_{mi}+x_{ki}$ is $1$-Lipschitz in $\mathbf{t}$ (with respect to $d_{\mathrm{tr}}$). Therefore $\delta_{j}$ is $N$-Lipschitz in $\mathbf{t}$. By the tropical descent rule:
$\delta_{j}(\mathbf{t}_{m+1})$ $\leq\delta_{j}(\mathbf{t}_{m})+Nd_{\mathrm{tr}}(\mathbf{t}_{m},\mathbf{t}_{m+1})$
@@
We have assumed that for each $m$, $\nabla f\neq\mathbf{0}$ and so there is some $i$ such that $\nabla f_{i}<0$, and by Lemma 25, there is some $k$ such that $i\in\text{argmin}_{\ell}(t_{m\ell}-x_{k\ell})$. Hence:
$\forall\,\ell:\quad t_{mj}-x_{kj}-t_{mi}+x_{ki}$ $\geq t_{mj}-x_{kj}-t_{m\ell}+x_{k\ell},$
$\Rightarrow\quad t_{mj}-x_{kj}-t_{mi}+x_{ki}$ $\geq\delta_{j}(\mathbf{t}_{m})/N,$
@@
Finally, from the claims above, we conclude that $\forall m\geq M_{2},{i\leq N}$, we have $\delta_{i}(\mathbf{t}_{m})\leq\epsilon$, and so by our first claim, we have $d_{{\rm tr}}(\mathbf{t}_{m},V)\leq\epsilon$. ∎
As proven by comuaneci2024location, tropical location problems have minima in the tropical convex hull of the data. The theorem above says that according to tropical descent, only such minima can be stable. We cannot do much better in terms of convergence results for general location problems; they can have disconnected sub-level sets so we would not expect gradient methods to find a global minimum. We would require some degree of convexity to guarantee global solvability, which is exactly what the following corollary gives us; tropical descent will necessarily converge to a minimum of any $\triangle_{\min}$-star-quasi-convex function.
@@
Let $f(\mathbf{t})$ be a $\triangle_{\min}$-star-quasi-convex function, and let $(\mathbf{t}_{m})_{m\geq 1}$ be a sequence of points defined by tropical descent such that for all $m$: $\nabla f(\mathbf{t}_{m})\neq 0$. Then there is some global minimum $\mathbf{t}^{*}$ such that for all $m\geq M_{2}$, $d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\epsilon$ where $M_{2}$ is as in Theorem 29.
This corollary follows from considering the 1-sample case of tropical location problems; Corollary 34 establishes convergence to the kernel of a $\triangle_{\min}$-star-quasi-convex function, which must necessarily be a global minimum. This convergence result holds even if $f$ does not have a unique minimum — it may be that $f$ has multiple global minima but a unique kernel, or $f$ may have a full dimensional set of global minima upon which $\nabla f(\mathbf{t})=\mathbf{0}$ and tropical descent terminates.
@@
Suppose $f$ is a max-tropical location problem with respect to the dataset $X=\{\mathbf{x}_{1},\dots,\mathbf{x}_{K}\}$. Let $(\mathbf{t}_{m})_{m\geq 1}$ be a sequence of points defined via max-tropical descent directions, such that for all $m$: $\nabla f(\mathbf{t}_{m})\neq 0$. Let $V=\operatorname{tconv}_{\min}(\mathbf{x}_{1},\dots,\mathbf{x}_{K})$, and define
$\Delta^{\prime}_{i}(\mathbf{t})=\sum_{{k\leq K}}\sum_{{j\leq N}}[t_{j}-x_{kj}-t_{i}+x_{ki}]^{+}.$
@@
The results above indicate convergence to $V$ or $\mathbf{t}^{*}$ in the tropical topology of $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, but $\mathbf{t}_{m}$ cannot converge in the Euclidean topology of $\mathbb{R}^{N}$ due to the divergent partial sums of $a_{m}$.
We conclude by establishing a foundational bound for the convergence rate of tropical descent; we take step sizes given by $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$, and assume derivatives have bounded tropical norm as is the case for piecewise linear 1-Lipschitz objective functions (with finitely many linear pieces). In this case, we prove convergence at a rate of $O(1/\sqrt{m})$.
@@
Suppose $f$ is a min-tropical location problem as in Theorem 29. Assume that there is some $L\in\mathbb{N}$ such that for all $\mathbf{t}_{m}$, $1/L\leq\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}\leq 2$, and assume further there is some $D$ such that $\forall m\in\mathbb{N}:\,\sum_{k\leq K}d_{\triangle_{\max}}(\mathbf{x}_{k},\mathbf{t}_{m})\leq KD$. We define $a_{m}=\alpha m^{-1/2}\|\nabla f(\mathbf{t}_{m})\|_{{\rm tr}}$. Then for all $m$ satisfying
${m\geq\max\left\{\left(\sqrt{2}+\frac{KDL}{2\alpha}-\sqrt{LKN-1}\right)^{2}-1,\left(\frac{2\alpha N}{D}\right)^{2}\right\},}$ (2)
@@
$d_{{\rm tr}}(\mathbf{t}_{m},V)\leq{\frac{2\sqrt{2}\alpha N}{\sqrt{m+1}-KDL/2\alpha+\sqrt{LKN-1}-\sqrt{2}}}=\epsilon_{m}.$